a car strats from rest has an accelerarion a=√t. a truck is moving with a uniform velocity of 4m/sec . at what distance will the car overtake the truck ?

Grade:12

1 Answers

Vikas TU
7 years ago
Already answered this qstn.
pls do check your previous one,
Again reposting,
 
For the car integerating two times to get the displacement eqn. as:
dv/dt = root(t)
v = 2t^(3/2)/3
and
dx/dt = 2t^(3/2)/3
x = 4t^(5/2)/15
For the truck,
x = 4t
solving both eqn.s we get,
t = 225^(1/3)
for the distance put t in x we get as:
x = 4*(225)^(1/3)